# Data-driven tool wear prediction in milling, based on a process-integrated single-sensor approach

### v2

@@ Abstract
Accurate tool wear prediction is essential for maintaining productivity and minimizing costs in machining. However, the complex nature of the tool wear process poses significant challenges to achieving reliable predictions. This study explores data-driven methods, in particular deep learning, for tool wear prediction. Traditional data-driven approaches often focus on a single process, relying on multi-sensor setups and extensive data generation, which limits generalization to new settings. Moreover, multi-sensor integration is often impractical in industrial environments. To address these limitations, this research investigates the transferability of predictive models using minimal training data, validated across two processes. Furthermore, it uses a simple setup with a single acceleration sensor to establish a low-cost data generation approach that facilitates the generalization of models to other processes via transfer learning. The study evaluates several machine learning models, including convolutional neural networks (CNN), long short-term memory networks (LSTM), support vector machines (SVM) and decision trees, trained on different input formats such as feature vectors and short-time Fourier transform (STFT). The performance of the models is evaluated on different amounts of training data, including scenarios with significantly reduced datasets, providing insight into their effectiveness under constrained data conditions. The results demonstrate the potential of specific models and configurations for effective tool wear prediction, contributing to the development of more adaptable and efficient predictive maintenance strategies in machining. Notably, the ConvNeXt model has an exceptional performance, achieving an 99.1% accuracy in identifying tool wear using data from only four milling tools operated until they are worn.
@@
Tool Wear Prediction, Signal Analysis, Machine Learning, Deep Learning, Milling, Vibration, Tool Condition Monitoring, Accelerometer.
@@ I Introduction
Economically, optimizing tool change times and extending tool life are essential to minimize downtime and maximize output, reducing overall operating costs. A primary concern in this domain is tool wear, which represents a challenge within the machining industry. At present, tool wear is primarily based on supplier recommendations and manual expertise [1]. However, because numerous factors influence tool wear during machining, its behavior is highly complex and variable, making it difficult to predict accurately using conventional or heuristic methods. So tool wear still occurs, often leading to disruptions in the manufacturing process. Ineffective management of tool wear, whether through premature or delayed identification, can result in a number of unwanted consequences, including unnecessary replacement costs and poor product quality [2]. Therefore, comprehensive research and development of intelligent tool condition monitoring (TCM) systems [3, 4, 5] is essential. TCM systems estimate the tool condition by passing sensor inputs through defined models, assisting machine operators in optimizing their work practices. The tool condition is characterized either through categorical classification, indicating distinct wear states, or through regression to estimate the wear condition. For regression, flank wear is usually applied as a measured variable, which is attempted to be predicted.
TCM can operate offline or online [6]:
  • Offline TCM, predictions are made post-process using additional data sources, which may include direct measurement methods. Tool wear is only identified with a significant delay or the process must be temporary interrupted.
@@
Indirect measurement methods of tool wear, such as force/torque, acceleration, temperature, motor currents, and acoustic sensors [7], produce data that must be processed and analyzed using intelligent models to gain interpretable insights. In contrast, direct measurement techniques, including image recognition, optical measurements and durability assessments, offer more straightforward data interpretation. The indirect method, or online TCM, requires no downtime and is therefore the preferred approach in many studies. Research in this area typically falls into two categories: physic-based models [8] and data-driven models [9]. Physic-based models offer detailed theoretical explanations and mechanism analysis, such as monitoring force modeling coefficients during cutting to track tool wear. However, due to the complex non-linear nature of cutting processes, it is difficult to parameterize and some effects, such as cutting temperature or lubrication conditions, are ignored to simplify the model. This limits the accuracy of the model when there is no clear physical mechanism. As a subset of Industry 4.0, rapid advances in computing, digitization and artificial intelligence (AI) have made it easier to implement data-driven models. These models use deep learning techniques, a subset of machine learning, to correlate sensor readings with tool wear, eliminating the need for a deep understanding of the physical processes. This enables faster implementation of TCM’s and makes data-driven approaches more accessible to diverse machining configurations, which is why this study used a data-driven method.
@@ II Related Work
To provide context for the single-sensor approach, we still introduce several multi-sensor systems (MSS) as a basis for comparison. A multi-sensor system has the advantage of combining diverse sensor types and thus different signal inputs. In addition, the same sensor type can record data at different positions, typically near the tool, the workpiece, or the spindle to capture specific aspects of the milling process. One approach in multi-sensor systems is to focus on a single area, similar seen in simpler sensor setups. In some studies, sensor placement focuses solely on the spindle, measuring spindle vibrations across multiple axes, spindle current, spindle force and the machine’s axis positions [10]. Other studies focus on the workpiece, using a stationary dynamometer under the workpiece (a ”force plate”) with additional sensors for vibration and spindle current monitoring in different positions [11].
However, most papers employ a multi-sensor approach by distributing sensors at various locations. These setups often use a force plate as the primary data source, supplemented by additional sensors for vibration and current measurements positioned elsewhere on the machine [12]. With multi-sensor setups, the data volume is inherently larger and can be further enhanced by adding a few hand-crafted features for training data-driven models. Various model architectures have been applied in MSS research, including random forest [13], support vector machines [14], simple neural networks [15] and convolutional neural networks (CNN) [16, 17, 18]. Among these, long short-term memory (LSTM) networks are frequently used due to their ability to handle time-series data [19, 20]. A common limitation across these MSS-based studies is the invasive machine modifications needed to install sensors, which often require specialized setups and time-consuming sensor installation and removal. Consequently, these studies are typically limited to testing on a single machine or using publicly available datasets [21, 22, 23] that were originally employed in a competitive context. Most public datasets still consist of data from only one machine.
@@
Overall, these studies demonstrate promising results within their specific use cases, showing that models can identify correlations between sensor measurements and tool wear. However, model validation has typically been limited to a single milling process, with only variations in the process parameters. As a result, while the models were proven capable of capturing wear-related patterns in this context, their ability to generalize or identify universal patterns in wear behavior, remains largely unexplored. Another limitation is that most studies conducted milling processes in a dry state to accelerate wear. In industrial applications, milling is often performed with a cooling lubricant, which affects tool wear differently. This discrepancy between experimental conditions and real-world practices makes it challenging to assess how well these research findings would translate to most industry applications.
In this study, we will address the limitations observed in previous research. As a result, this study makes the following contributions:
@@
  • Utilization of a single accelerometer sensor, which is straightforward to integrate into a range of machines and processes
@@
  • Comparison of multiple model architectures, including a ConvNeXt model, LSTM, decision tree and supper vector classifier (SVC).
@@ III-A Problem description and Pipeline
Data-driven models require large datasets to capture the non-linear relationships and random variations inherent in complex machine processes. In addition, these models are often tailored to the specific processes they are trained on. In a real world applications, however, processes can frequently change, and for companies to justify the investment in wear detection systems, these models need to be adaptable to new or evolving processes. The complexity of tool wear makes it challenging to develop a generalizable data-driven model that can apply across various processes, limiting the practical implementation of such research. To address this issue, this paper proposes a different approach by comparing various data-driven models and evaluating their performance based on the quantity of available training data. The training data is based on a single acceleration sensor, enabling fast and cost-effective model training for any process. It also makes it easier to adapt the model to new processes, making it more suitable for transfer learning. The training datasets contain acceleration data of vibrations during the milling process. From these raw signal transformed entities, relevant features are extracted with the objective of predicting tool wear. The problem is framed as a binary classification task, with the aim of determining if wear with a negative impact is present.
@@ III-B Data preprocessing
The accelerometer generates a continuous stream of acceleration data $\ddot{x}$ that includes not only the critical process information related to tool wear, but also all the surrounding recordings, such as spindle run-up, tool change and the transitional moments when the tool is starting or finishing its milling operation. To filter out these unwanted data, the signal is automatically processed by analyzing its peak-to-peak ($\ddot{x}_{\text{p2p}_{t}}$) over time steps t with a moving window of length w.
@@
A threshold is calculated to isolate the segments of the signal that correspond to the actual milling process. For this, the $\ddot{x}_{\text{p2p}_{t}}$ curve is sorted in ascending order ($\uparrow$) and then the average lowest value $m_{\text{p2p}}$ is determined as described below. Here N is the length of the $\ddot{x}_{\text{p2p}_{t}}$ curve.
@@
The $\alpha_{0}-\alpha_{2}$ represent empirically determined values: $\alpha_{0}=0.01$, $\alpha_{1}=0.03$, $\alpha_{2}=10$. The signal is also trimmed according to the duration for which it is below or above the threshold $th$, thereby preventing the misclassification of brief anomalies with comparable vibration energy. The data preprocessing is explained in Fig. 1.
@@ III-C Feature engineering
The acceleration data takes a measurement every 65 $\mu s$, giving a sampling rate of approximately 15,38 kHz. This leads to a Nyquist border with 7,89 kHz, so the data theoretically contains a large amount of information. Nevertheless, the data is derived from only a one-dimensional acceleration signal, posing a challenge in identifying the complete range of characteristics and correlations related to wear. To address this, several techniques have been developed to improve the interpretability of the signal for the model, often by reducing the amount of data while preserving the essential information. One such approach is signal transformation, where the signal can be analyzed in different domains: time domain, frequency domain, or time-frequency domain. In this paper, the Fourier or Welch transformation is used to transfer the signal in the frequency domain and the STFT for the time-frequency domain. As a result, the original one-dimensional signal is transformed into a multi-dimensional input for the model. Fig. 2a shows the different data transformations on the left. Optionally, the input can be further adjusted to reduce the input size. This is done by extracting features. Popular features are typical statistical features such as mean, peak-to-peak, kurtosis, and skewness. However, there are also special or self-developed features that are used in this paper. The most important features are listed in Table I.
@@ III-D Different models
Fig. 2a illustrates the structure of the models. First, there is the input, followed by the type of feature extraction (i, ii) and the classifier used (1, 2, 3, 4). The first type of feature extraction (i) is a model backbone. In this work, an advanced and modern CNN architecture, the ConvNeXt architecture [30], is selected as the backbone. This model builds upon the successful ResNet framework [31] and incorporates enhancements inspired by modern model architectures like the Swin Transformer[32]. Benchmark studies have demonstrated that this model is at least as effective as the Vision Transformer [33]. In the Model ”ConvNeXt”, wear is predicted from the extracted features of the STFT with fully connected layers (1). In the model ”ConvNeXtLSTM”, a sequence of features is extracted from a sequence of STFT’s and processed with LSTM’s to subsequently predict wear with fully connected layers as well (2).
In the second type of feature extraction (ii), features are derived from the acceleration data using proprietary algorithms, producing a feature vector of 20 selected features as input. This feature vector is utilized by the model ”SVC” with a Support Vector classifier (3) and by the model ”DTreeC” with a decision tree classifier (4). Fig. 2b shows the different models in detail.
Figure 2: (a) The diagram illustrates the transformation of acceleration data into different domains and the subsequently extraction of features using either a CNN architecture (i) or proprietary algorithms (ii). These features are then utilized by various classifiers (1, 2, 3, 4) to predict wear. (b) The architectures of the four classifiers used in this paper, categorized by their input type, are shown. Images from the following sources were used to create this diagram: [34, 35]

### v3

@@ Abstract
Accurate tool wear prediction is essential for maintaining productivity and minimizing costs in machining. However, the complex nature of the tool wear process poses significant challenges to achieving reliable predictions. This study explores data-driven methods, in particular deep learning, for tool wear prediction. Traditional data-driven approaches often focus on a single process, relying on multi-sensor setups and extensive data generation, which limits generalization to new settings. Moreover, multi-sensor integration is often impractical in industrial environments. To address these limitations, this research investigates the transferability of predictive models using minimal training data, validated across two processes. Furthermore, it uses a simple setup with a single acceleration sensor to establish a low-cost data generation approach that facilitates the generalization of models to other processes via transfer learning. The study evaluates several machine learning models, including transformer-inspired convolutional neural networks (CNN), long short-term memory networks (LSTM), support vector machines (SVM), and decision trees, trained on different input formats such as feature vectors and short-time Fourier transform (STFT). The performance of the models is evaluated on two machines and on different amounts of training data, including scenarios with significantly reduced datasets, providing insight into their effectiveness under constrained data conditions. The results demonstrate the potential of specific models and configurations for effective tool wear prediction, contributing to the development of more adaptable and efficient predictive maintenance strategies in machining. Notably, the ConvNeXt model has an exceptional performance, achieving 99.1% accuracy in identifying tool wear using data from only four milling tools operated until they are worn.
@@
Tool wear prediction, Machine learning, Milling, Industry-ready models, Data-driven approach, Accelerometer
@@ I Introduction
Economically, optimizing tool change times and extending tool life are essential to minimize downtime and maximize output, reducing overall operating costs. A primary concern in this domain is tool wear, which represents a challenge within the machining industry. At present, tool wear is primarily based on supplier recommendations and manual expertise [1]. However, because numerous factors influence tool wear during machining, its behavior is highly complex and variable, making it difficult to predict accurately using conventional or heuristic methods. So, tool wear still occurs, often leading to disruptions in the manufacturing process. Ineffective management of tool wear, whether through premature or delayed identification, can result in a number of unwanted consequences, including unnecessary replacement costs and poor product quality [2]. Therefore, comprehensive research and development of intelligent tool condition monitoring (TCM) systems [3, 4, 5] is essential. TCM systems estimate the tool condition by passing sensor inputs through defined models, assisting machine operators in optimizing their work practices. The tool condition is characterized either through categorical classification, indicating distinct wear states, or through regression to estimate the wear condition. For regression, flank wear is usually applied as a measured variable, which is attempted to be predicted. TCM can operate offline or online [6]:
  • Offline TCM, predictions are made post-process using additional data sources, which may include direct measurement methods. Tool wear is only identified with a significant delay or the process must be temporarily interrupted.
@@
Indirect measurement methods of tool wear, such as force/torque, acceleration, temperature, motor currents, and acoustic sensors [7], produce data that must be processed and analyzed using intelligent models to gain interpretable insights. In contrast, direct measurement techniques, including image recognition, optical measurements, and durability assessments, offer more straightforward data interpretation. The indirect method, or online TCM, requires no downtime and is therefore the preferred approach in many studies. Research in this area typically falls into two categories: physic-based models [8] and data-driven models [9]. Physic-based models offer detailed theoretical explanations and mechanism analysis, such as monitoring force modeling coefficients during cutting to track tool wear. However, due to the complex non-linear nature of cutting processes, it is difficult to parameterize and some effects, such as cutting temperature or lubrication conditions, are ignored to simplify the model. This limits the accuracy of the model when there is no clear physical mechanism. As a subset of Industry 4.0, rapid advances in computing, digitization, and artificial intelligence (AI) have made it easier to implement data-driven models. These models use deep learning techniques, a subset of machine learning, to correlate sensor readings with tool wear, eliminating the need for a deep understanding of the physical processes. However, a fundamental understanding of sensor selection and input types remains necessary for an effective data-driven approach to tool wear prediction. Fig. 1 presents an overview of this essential knowledge and the sensors most commonly used for both direct and indirect measurement methods, with this study focusing on indirect measurements.
Figure 1: The influence of tool wear and the sensors that can be employed for its monitoring.
The Fig. 1 demonstrates how various types of wear, such as flank wear, lead to changes in the tool geometry (Fig. 1, (1)), which in turn affect the cutting forces. These forces can be categorized into static (mean) and dynamic components. As cutting forces directly interact with wear, they serve as primary indicators for TCM’s. In milling, dynamic cutting forces play a particularly important role, as the intermittent engagement of cutting edges results in complex time-dependent interactions that can be effectively analyzed in the frequency domain. Direct measurement of cutting forces is possible through the use of a dynamometer, which is a widely used sensor in research. Nevertheless, the high cost and invasive installation of dynamometers limit their industrial application. A popular alternative are current sensors, which are often already integrated into machines. These sensors provide an indirect way of force measurement by correlating current consumption with the required cutting force. In a closed-loop system, dynamic forces are a contributing factor to vibrations, which are induced based on the process and machine dynamics (Fig. 1, (2)). These vibrations result in an uneven machined surface (Fig. 1, (3)), which subsequently influences the cutting thickness in the following tool engagement (Fig. 1, (4)). Consequently, the cutting thickness varies over time, further affecting the dynamic cutting forces. In this closed-loop system, the only indirect measurement method is vibrations, which can be measured with accelerometers. Accelerometers offer several advantages, including cost-effectiveness, minimal calibration requirements, and seamless integration into machines. The high sampling rates of accelerometers facilitate detailed time-domain analysis, capturing transient effects such as short peaks resulting from crater wear on the cutting edge. Moreover, accelerometers are particularly adept at frequency-domain analysis due to their capacity to cover a broad frequency range and their ability to be strategically positioned to minimize ambient noise. The utilization of advanced signal processing techniques, such as the Fast Fourier Transform (FFT), wavelet analysis, or the Welch method, ensures effective filtration of frequency background noise. In summary, accelerometers enable efficient and effective implementation of TCM’s, which is why this study employs a data-driven approach using accelerometers.
@@ II Related Work
To provide context for the single-sensor approach, we still introduce several multi-sensor systems (MSS) as a basis for comparison. A multi-sensor system has the advantage of combining diverse sensor types and thus different signal inputs. In addition, the same sensor type can record data at different positions, typically near the tool, the workpiece, or the spindle to capture specific aspects of the milling process. One approach in multi-sensor systems is to focus on a single area, similar seen in simpler sensor setups. In some studies, sensor placement focuses solely on the spindle, measuring spindle vibrations across multiple axes, spindle current, spindle force, and machine axis positions [10]. Other studies focus on the workpiece, using a stationary dynamometer under the workpiece (a ”force plate”) with additional sensors for the monitoring of spindle current and vibration [11].
However, most papers employ a multi-sensor approach by distributing sensors at various locations. These setups often use a force plate as the primary data source, supplemented by additional sensors for vibration and current measurements positioned elsewhere on the machine [12]. With multi-sensor setups, the data volume is inherently larger and can be further enhanced by adding a few handcrafted features for training data-driven models. Various model architectures have been applied in MSS research, including random forest [13], support vector machines [14], simple neural networks [15], and convolutional neural networks (CNN) [16, 17, 18]. Among these, long short-term memory (LSTM) networks are frequently used due to their ability to handle time-series data [19, 20]. A common limitation across these MSS-based studies is the invasive machine modifications needed to install sensors, which often require specialized setups and time-consuming sensor installation and removal. Consequently, these studies are typically limited to testing on a single machine or using publicly available datasets [21, 22, 23] that were originally employed in a competitive context. Most public datasets still consist of data from only one machine.
@@
Overall, these studies demonstrate promising results within their specific use cases, showing that models can identify correlations between sensor measurements and tool wear. However, model validation has typically been limited to a single milling process, with only variations in the process parameters. As a result, while the models were proven capable of capturing wear-related patterns in this context, their ability to generalize or identify universal patterns in wear behavior, remains largely unexplored. Another limitation is that most studies conducted milling processes in a dry state to accelerate wear. In industrial applications, milling is often performed with a cooling lubricant, which affects tool wear differently. This discrepancy between experimental conditions and real-world practices makes it challenging to assess how well these research findings would translate to most industry applications. In this study, we will address the limitations observed in previous research. As a result, this study makes the following contributions:
@@
  • Utilization of a single accelerometer sensor, which is straightforward to integrate into a range of machines and processes.
@@
  • Performance analyisis of multiple model architectures, including a ConvNeXt model, LSTM, decision tree, and supper vector classifier (SVC).
@@ III-A Problem description and Pipeline
Data-driven models require large datasets to capture the non-linear relationships and random variations inherent in complex machining processes. In addition, these models are often tailored to the specific processes they are trained on. In real world applications, however, processes can frequently change, and for companies to justify the investment in wear detection systems, these models need to be adaptable to new or evolving processes. The complexity of tool wear makes it challenging to develop a generalizable data-driven model that can be applied across various processes, limiting the practical implementation of such research. To address this issue, this paper proposes a different approach by comparing various data-driven models and evaluating their performance based on the quantity of available training data. The training data is based on a single acceleration sensor, enabling fast and cost-effective model training for any process. It also makes it easier to adapt the model to new processes, making it more suitable for transfer learning. The training datasets contain acceleration data of vibrations during the milling process. From these raw signal transformed entities, relevant features are extracted with the objective of predicting tool wear. The problem is framed as a binary classification task with the aim of determining whether wear with a negative impact is present. This simplification aligns with industry needs, as companies primarily require clear indications of when wear starts to degrade machining quality, rather than a continuous regression-based wear estimation.
@@ III-B Data preprocessing
The accelerometer generates a continuous stream of acceleration data $\ddot{x}$ that includes not only the critical process information related to tool wear, but also all the surrounding recordings, such as spindle run-up, tool change, and the transitional moments when the tool is starting or finishing its milling operation. To filter out these unwanted data, the signal is automatically processed by analyzing its peak-to-peak ($\ddot{x}_{\text{p2p}_{t}}$) over time steps t with a moving window of length w.
@@
A threshold is calculated to isolate the segments of the signal that correspond to the actual milling process. For this, the $\ddot{x}_{\text{p2p}_{t}}$ graph is sorted in ascending order ($\uparrow$) and then the average lowest value $m_{\text{p2p}}$ is determined as described below. Here N is the length of the $\ddot{x}_{\text{p2p}_{t}}$ curve.
@@
The $\alpha_{i}$ represent empirically determined values, with $\alpha_{0}=0.01$, $\alpha_{1}=0.03$, $\alpha_{2}=10$. The signal is also trimmed according to the duration for which it is below or above the threshold $th$, thereby preventing the misclassification of brief anomalies with comparable vibration energy. The data preprocessing is explained in Fig. 2.
@@ III-C Feature engineering
The acceleration data takes a measurement every 65 $\mu s$, giving a sampling rate of approximately 15,38 kHz. This leads to a Nyquist border with 7,89 kHz, so the data is highly frequency sensitive and prone to short time-variant fluctuations or changes. But one minute of data already contains almost a million data points packed all in a one-dimensional acceleration signal. This complexity posing a challenge in identifying the complete range of characteristics and correlations related to wear. To reduce the complexity current approaches are data transformations and/or feature extraction. These approaches often reducing the amount of data while preserving the essential information.
Data or signal transformations are essential to the analysis of vibration signals, particularly in the frequency and time-frequency domains, as frequency characteristics provide crucial insight into the milling process. Traditional techniques such as the Fast Fourier Transform (FFT) are still widely used. But in milling, how frequencies change over time is also important, making time-frequency transformations more relevant. Common methods in this area include the Wavelet Transform, Hilbert Huang Transform, Empirical Mode Decomposition (EMD) and the Short-Time Fourier Transform (STFT). In this study, the STFT is used in combination with the Welch method to transform the data. The STFT provides time-resolved frequency analysis by applying the Fourier transform to short, overlapping segments of the signal. The Welch method further improves spectral estimation by averaging segments of the Fourier transform, improving frequency accuracy and reducing noise. The resulting transformation produces a two-dimensional, picture-like representation where the y-axis represents frequency $f$ and the x-axis represents time $t$. The formula for an STFT is as follows
$S[f,t]=\sum_{m=0}^{M-1}\ddot{x}_{t}[m]e^{-2j\pi f\frac{m}{M}}$ (4)
In the STFT, $\ddot{x}_{t}[m]$ represents a windowed segment of the signal to which the Fourier transform is applied. A fundamental property of the STFT is its fixed frequency distribution for every application. This ensures that, irrespective of the signal’s actual frequency content, the energy components are consistently mapped to the same frequency bins in the STFT representation. A further advantage of the STFT is that it does not increase the overall data volume when transforming a one-dimensional signal into a two-dimensional representation, which makes it computationally efficient, particularly when using convolution-based algorithms, which benefit from the structured nature of the transformed data. In contrast, other time-frequency transformations, such as the wavelet transform and empirical mode decomposition (EMD), introduce a significantly higher computational burden. The wavelet transform necessitates repeated convolutions with different basis functions, resulting in increased processing time, while EMD involves an iterative decomposition process that is inherently computationally expensive. These methods yield high-dimensional outputs that require substantial resources for further processing and possess high intrinsic computational costs, rendering them challenging to implement in real-time industrial applications. To further reduce noise and data complexity, this study employs the Welch method in combination with the STFT. Fig. 3a illustrates the different data transformation approaches, with the corresponding transformations shown on the left. The transformed data can now be used as input for a data-driven model, allowing it to perform feature extraction automatically. Alternatively, manually engineered features can be generated using custom algorithms to further reduce the input dimensionality before feeding the data into a classification model. In this study, both approaches are explored.
For the purpose of wear detection in milling and vibration signal analysis, well-established features have been selected. Statistical features remain fundamental and include metrics such as mean, peak-to-peak amplitude, kurtosis, and skewness. In addition to these, more specialized features are considered to capture changes in signal complexity, which often indicate wear. In the frequency domain, an increase in signal complexity is characterized by the presence of multiple frequency components, which is quantified using the Spectral Entropy feature, which measures how evenly energy is distributed across different frequencies. In the time domain, disorder can be assessed using the Fractal Dimension, which describes the irregularity of vibration patterns. Another important feature extraction method involves separating the signal’s energy in the frequency domain into periodic and aperiodic components. An increase in aperiodic energy, i.e. energy from frequencies not associated with fundamental process frequencies, such as the tooth entry frequency, can be an indicator of tool wear. The features employed in this study, along with their respective descriptions, are outlined in Tab. I, II.
@@ III-D Different models
One objective of this research is to identify the most suitable model for an industry-ready single-sensor installation for tool wear detection. To this end, a range of state-of-the-art architectures are evaluated, differing in complexity and computational cost. The aim is to determine the level of model intelligence required for effective wear prediction. Fig. 3a illustrates the structures underlying the different models. The modeling process consists of two key components: feature extraction from the signal (Fig. 3, (i, ii)) and classification (Fig. 3, (1,2,3,4)). The first feature extraction approach (Fig. 3, (i)) relies on a model-based backbone, which is computationally more intensive but allows the model to learn relevant features autonomously. As discussed in the previous chapter, the STFT is used as the input representation. The task is then framed as an image classification problem, where the model must recognize wear conditions from the transformed data. At present, two leading architectures dominate image classification: Convolutional Neural Networks (CNN) [30, 31] and Vision Transformers (ViTs)[32]. While ViTs have demonstrated strong performance in the field of computer vision, CNNs remain superior in many domains, particularly when training data is limited and computational efficiency is crucial [33]. Recent advances in deep learning have led to hybrid CNN-ViT architectures [34], combining the strengths of both approaches [35]. Among these, CNN-based models continue to be the dominant component. In this study, the CNN-ViT hybrid model, ConvNeXt, is selected as the backbone. ConvNeXt builds on the widely used ResNet framework [36], incorporating architectural enhancements inspired by the Swin Transformer [37]. Benchmark studies show that ConvNeXt performs on par with or better than Vision Transformer in many cases [38].
As demonstrated in Fig. 3b, the first two models utilize ConvNeXt as the backbone.
  1. 1.
    Model ”ConvNeXt”: Employs a simple fully connected layer as the classifier. While this model is computationally efficient, it processes STFTs independently and is unable to account for temporal dependencies beyond the information embedded in each individual STFT.
  2. 2.
    Model ”ConvNeXtLSTM”: Designed to handle sequential wear data, and processes a sequence of STFTs. Each STFT is first passed through the ConvNeXt backbone, generating a sequence of features. These features are then processed by an LSTM-based classifier with three layers (64, 32, and 32 LSTM units, respectively), followed by a fully connected layer. To mitigate overfitting, dropout (10%) is applied to the first two LSTM layers.
The third and fourth models rely on manually engineered features (Fig. 3b, (ii)). In this approach, a moving window is employed to extract 20 handcrafted features at each step. These extracted features are subsequently fed into classical machine learning classifiers:
  1. 3.
    Model ”SVC”: A Support Vector Classifier with a radial basis function (RBF) kernel.
  2. 4.
    Model ”DTreeC”: A Decision Tree Classifier using the Gini impurity criterion, without depth limitation.
Figure 3: (a) The diagram illustrates the transformation of acceleration data into different domains and the subsequently extraction of features using either a CNN architecture (i) or proprietary algorithms (ii). These features are then utilized by various classifiers (1, 2, 3, 4) to predict wear. (b) The architectures of the four classifiers used in this paper, categorised by their input type, are shown. Images from the following sources were used to create this diagram: [39, 40]